A rectangular parcel of land is 50ft wide. The length of a diagonal between opposite corners is 10ft more than the length of the parcel. What is the length of the parcel?
Answer by checkley71(8403)

L=LENGTH
DIAGONAL=(L+10)
W=50
USING THE PYTHAGOREON THEROUM A^2+B^2=C^2
L^2+50^2=(L+10)^2
L^2+2500=L^2+20L+100
20L=2500-100
20L=2400
L=2400/20
L=120 IS THE LENGTH.
120+10=130 IS THE DIAGONAL.
PROOF
120^2+50^2=130^2
14400+2500=16900
16900=16900